# Cooling of quark stars from perturbative QCD

### v1

@@ I Introduction
Neutron stars are believed to occupy a region of the Quantum Chromodynamics (QCD) phase diagram characterized by low temperatures and high baryon densities. While it may be surprising at first glance to call neutron stars cold, since they are born in supernova explosions with temperatures of $T\sim 10^{11}\ \mathrm{K}\sim 10\ \mathrm{MeV}$ and cool over time to temperatures in the keV range Schmitt (2010); Burgio et al. (2021), these values are small compared to the energy scale set by QCD. For example, the temperature of the deconfinement transition at vanishing chemical potential is around $T_{c}\sim(150-160)\ \mathrm{MeV}$ Cheng et al. (2010); Borsanyi et al. (2010, 2014); Kanaya et al. (2018), and the quark chemical potential to be found in the interior of neutron stars is typically within the range of hundreds of MeV Schaffner-Bielich (2020). Because neutron star temperatures are far below such scales, their matter is usually treated within the framework of cold and dense QCD Schaffner-Bielich (2020); Lovato et al. (2022).
@@
In this work, we investigate the thermal evolution of quark stars – hypothetical compact objects composed entirely of deconfined strange quark matter Bodmer (1971); Witten (1984); Terazawa (1989) – with and without a hadronic crust, using an equation of state derived from perturbative QCD to order $\alpha_{s}$. We incorporate the renormalization group running of the strong coupling, $\alpha_{s}$, and the strange quark mass, $m_{s}$. We focus on the thermal emission of these objects. As will be discussed in detail, such stars lose thermal energy via two primary mechanisms: neutrino emission from the bulk and photon emission from the surface Tsuruta (1998). During the early stages of evolution, neutrino emission dominates the cooling and is highly sensitive to the microscopic composition of the star. Among the various neutrino processes, the most efficient is the direct Urca process Gamow and Schoenberg (1940, 1941). In nucleonic matter, this process is only kinematically allowed when the proton fraction exceeds a critical threshold Lattimer et al. (1991); Schaab et al. (1999). In quark matter, however, the direct Urca process can occur if interactions between quarks are included (i.e., if the strong coupling $\alpha_{s}$ is non-zero) and if electrons are present Blaschke et al. (2000).
@@ II.1 Evolution equations
$\dfrac{\partial T}{\partial t}=-\dfrac{1}{4\pi r^{2}e^{\Phi}C_{\nu}}\sqrt{1-\dfrac{2Gm}{c^{2}r}}\dfrac{\partial}{\partial r}(e^{2\Phi}L)-\dfrac{\epsilon_{\nu}}{C_{\nu}}e^{\Phi}\,$, (1)
@@
$L=-4\pi\kappa r^{2}\sqrt{1-\dfrac{2Gm}{c^{2}r}}e^{-\Phi}\dfrac{\partial}{\partial r}(Te^{\Phi})\,$, (2)
@@ III.1 Quark matter
$d\rightarrow u+e+\bar{\nu}_{e}\,$, $u+e\rightarrow d+\nu_{e}\,$, (6)
$s\rightarrow u+e+\bar{\nu}_{e}\,$, $u+e\rightarrow s+\nu_{e}\,$,
$s+u\leftrightarrow d+u\,$,
@@ III.2 Crust
$P(\epsilon)=\begin{cases}P_{\mathrm{BPS}}(\epsilon),&\text{if }P<P_{\mathrm{crust}}\\ P_{\mathrm{pQCD}}(\epsilon),&\text{if }P\geq P_{\mathrm{crust}}.\end{cases}$ (8)
@@ IV Results
In what follows, we compute cooling curves, i.e., the evolution of $L_{\infty}$ with stellar age, for quark stars with and without a crust, considering configurations with mass $M=1.4M_{\odot}$. The results obtained with the pQCD equation of state are shown in gray, where the cases with $\bar{\Lambda}=2\mu$ (dashed line) and $\bar{\Lambda}=4\mu$ (solid line) define the lower and upper limits of the gray band. For comparison, we also compute the cooling curve using the MIT bag model with $\mathcal{O}(\alpha_{s})$ corrections Farhi and Jaffe (1984) in which the parameters are set to $\alpha_{s}=0.1$, $m_{s}=150\ \mathrm{MeV}$ and $B^{1/4}=140\ \mathrm{MeV}$. This result is shown in black in the cooling curves.
@@
In both cases, the initial temperature drop occurs in the outer layers, where thermal photon emission – used as a boundary condition in the cooling equations – takes place, while the core remains relatively hotter. The subtle difference in the initial conditions between the two limits arises from the redshift factor $e^{\Phi(r)}$, which differs between the configurations due to their different metric functions. By $t=10^{-3}$ yr, the temperature profiles are quite similar across both configurations, with some differences: the upper limit exhibits a slightly higher temperature throughout most of the stellar profile compared to the lower limit, on the order of $10^{7}\ \mathrm{K}$. Moreover, in contrast to the interior, the surface temperature is slightly lower by a comparable amount.
We observe a similar behavior at $t=10^{-1}$ yr, with a hotter interior and a steeper temperature gradient in the outer layers. However, at this point we begin to access the differences in the thermalization process between the limits considered. The configuration with $\bar{\Lambda}=4\mu$ appears to thermalize more rapidly than that with $\bar{\Lambda}=2\mu$, since it exhibits a more extended region within the star that is isothermal with the center. By the time both configurations reach $\mathrm{t}=1$ yr, the star becomes nearly isothermal, with a temperature difference between the center of the star ($r=0$) and the surface ($r=R_{\rm star}$) of the order of $10^{3}\ \mathrm{K}$ in both cases. This difference decreases further to the order of $10^{0}$ K at $t=8$ yr for $\bar{\Lambda}=2\mu$ and $t=13$ yr for $\bar{\Lambda}=4\mu$.
@@
For quark stars featuring a hadronic crust, we compute the thermal evolution using two distinct numerical approaches. The first method explicitly accounts for the presence of the gap between the quark core and the crust. In this case, when constructing the radial grid for the calculations, we place a grid point at the location of the gap by employing a sufficiently small radial step size ($dr$). The second method neglects the gap entirely, initiating the crustal grid at the neutron drip density and continuing outward, down to a density of $\rho=10^{10}~\mathrm{g\,cm^{-3}}$. We find that, while the resulting cooling band remains unchanged when comparing the two methods, the internal temperature profiles of the star evolve differently in early stages of thermal relaxation, as will become clear in the subsequent analysis. Notably, the first method, which resolves the gap explicitly, results in significantly increased computational time.
@@
Furthermore, the luminosity band for the crusted quark star is broader than that of the bare configuration and does not show a significant contraction at early times. Instead, the band gradually narrows as the system evolves toward $t\sim 10^{5}$ yr, yet remains broader than in the bare case. This behavior signals that the sensitivity to the choice of the renormalization scale becomes more pronounced in the cooling of quark stars with a hadronic crust.
@@
In Figure 8, we show the temperature profile evolution for the two modeling approaches for a quark star with a crust. The top panels correspond to the case where the thermal gap at the crust–core boundary is neglected, whereas the bottom panels show the results when this gap is taken into account. The most significant differences arise during the early evolutionary stages. At these times, as shown for $t=10^{-3}$ yr and $t=10^{-1}$ yr, the temperature profile exhibits a clear discontinuity between the crust (for densities $\rho\leq 4\times 10^{10}\ \mathrm{g}\ \mathrm{cm}^{-3}$) and the core ($\rho>4\times 10^{10}\ \mathrm{g}\ \mathrm{cm}^{-3}$) when we take the gap into account. The core rapidly becomes isothermal, while the crust retains higher temperatures, except in the outermost layers where surface photon emission plays an important role.
In contrast, when the thermal gap is not included, the transition between crust and core is smoother, and the core takes longer to reach an isothermal state. Despite these early-time differences in the internal temperature distribution, both models converge to nearly identical thermal profiles by $t=1$ yr. The differences in internal temperature evolution, however, do not translate into observable consequences: the resulting redshifted luminosity $L_{\infty}$ remains identical across the entire time span considered in this study.
Beyond the variations in the cooling band behavior, the temperature profile evolution further highlights differences between the bare and crusted quark star models. In the presence of a crust, a pronounced temperature gradient develops, with the two regions cooling at different rates. This behavior resembles that of neutron star cooling, where efficient neutrino emission in the core causes it to cool more rapidly than outer layers, leading to a “cold wave” propagating outward toward the crust. This process delays the onset of an isothermal profile. Specifically, the time required for the temperature difference between the center ($r=0$) and the boundary ($r=R_{\rm boundary}$) to decrease to a value on the order of $10^{3}$ K is approximately $240$ yr for $\bar{\Lambda}=2\mu$ and $700$ yr for $\bar{\Lambda}=4\mu$. Additionally, for quark stars with a crust, a temperature disparity persists between the two limits of the band. While in the bare case the temperature profiles for both limits converge to nearly identical values by $t=1$ yr, the crusted case maintains a residual difference between the core and the crust, which, although gradually reduced, remains present up to $t=10^{5}$ yr.
@@ V Summary and outlook
Motivated by these findings, we extended our analysis to quark stars with a hadronic crust. The motivation arises from the observation that the uncertainty associated with the pQCD EoS is more significant at early stages of thermal evolution, when the star has not yet achieved internal isothermality and is primarily cooling via neutrino emission. Given the known role of the crust as a thermal insulator, which delays the establishment of an isothermal interior, we investigated how the uncertainty from the pQCD EoS influences this configuration. As expected, the temperature evolution in the crusted case is less steep than in the bare case, bringing the model predictions to a good agreement with a portion of the observational data, as illustrated in Figure 7. However, in contrast to the bare case, the cooling band in the crusted configuration does not exhibit a significant narrowing throughout the timescale considered (up to $t=10^{5}$ yr), indicating a more persistent sensitivity to the EoS parameters.
We also computed the cooling curves obtained using the MIT bag model with $\mathcal{O}(\alpha_{s})$ corrections. For both bare and crusted quark stars, the MIT bag model results lie within the uncertainty band defined by the variation of $\bar{\Lambda}$ in the pQCD EoS. The only exception occurs at the very beginning of the evolution of the quark star with a crust, in which the MIT bag model curve lies outside the band. It enters the band before the star reaches $10^{-1}$ years.
It is well established that color superconductivity can modify both the neutrino emissivity and the heat capacity of quark matter Shovkovy and Ellis (2002); Blaschke et al. (2006); Noda et al. (2013); Schmitt and Shternin (2018). A natural extension of the present study would therefore be to investigate the cooling of quark stars in different color-superconducting phases, while consistently incorporating the uncertainties from the pQCD equation of state. Another relevant direction is to refine the microphysical inputs by implementing up-to-date calculations of neutrino emissivity in quark matter Suleiman et al. (2023); Sedrakian (2024); Alford et al. (2024).

### v2

@@ I Introduction
Neutron stars are believed to occupy a region of the Quantum Chromodynamics (QCD) phase diagram characterized by low temperatures and high baryon densities. While it may be surprising at first glance to call neutron stars cold, since they are born in supernova explosions with temperatures of $T\sim 10^{11}\ \mathrm{K}\sim 10\ \mathrm{MeV}$ and cool over time to temperatures in the keV range Schmitt (2010); Burgio et al. (2021), these values are small compared to the energy scale set by QCD. For example, the temperature of the deconfinement transition at vanishing chemical potential is around $T_{c}\sim(150-160)\ \mathrm{MeV}$ Cheng and others (2010); Borsanyi et al. (2010, 2014); Kanaya et al. (2018), and the quark chemical potential to be found in the interior of neutron stars is typically within the range of hundreds of MeV Schaffner-Bielich (2020). Because neutron star temperatures are far below such scales, their matter is usually treated within the framework of cold and dense QCD Schaffner-Bielich (2020); Lovato and others (2022).
@@
In this work, we investigate the thermal evolution of quark stars – hypothetical compact objects composed entirely of deconfined strange quark matter Bodmer (1971); Witten (1984); Terazawa (1989) – with and without a hadronic crust, using an equation of state derived from perturbative QCD to order $\alpha_{s}$. We incorporate the renormalization group running of the strong coupling, $\alpha_{s}$, and the strange quark mass, $m_{s}$. We focus on the thermal emission of these objects. As will be discussed in detail, such stars lose thermal energy via two primary mechanisms: neutrino emission from the bulk and photon emission from the surface Tsuruta (1998). During the early stages of evolution, neutrino emission dominates the cooling and is highly sensitive to the microscopic composition of the star. Among the various neutrino processes, the most efficient is the direct Urca process Gamow and Schoenberg (1940, 1941). In nucleonic matter, this process is only kinematically allowed when the proton fraction exceeds a critical threshold Lattimer et al. (1991); Schaab et al. (1999); Mendes et al. (2025). In quark matter, however, the direct Urca process can occur if interactions between quarks are included (i.e., if the strong coupling $\alpha_{s}$ is non-zero) and if electrons are present Blaschke et al. (2000).
@@ II.1 Evolution equations
[…]
@@
[…]
@@ III.1 Quark matter
$aligned&d\rightarrow u+e+\bar{\nu}_{e}\,,&&u+e\rightarrow d+\nu_{e}\,,\\ &s\rightarrow u+e+\bar{\nu}_{e}\,,&&u+e\rightarrow s+\nu_{e}\,,\\ &&s+u\leftrightarrow d+u\,$, (6)
@@ III.2 Crust
$P(\epsilon)=\cases{P}_{\mathrm{BPS}}(\epsilon),&\text{if }P<P_{\mathrm{crust}}\\ P_{\mathrm{pQCD}}(\epsilon),&\text{if }P\geq P_{\mathrm{crust}}.$ (8)
@@ IV Results
[…]
@@
The subtle difference in the initial conditions between the two limits arises from the redshift factor $e^{\Phi(r)}$, which differs between the configurations due to their different metric functions. In both cases, the initial temperature drop occurs in the outer layers, where thermal photon emission – used as a boundary condition in the cooling equations – takes place, while the core remains relatively hotter. By $t=10^{-3}$ yr, the temperature profiles are quite similar across both configurations, with some differences: the upper limit exhibits a slightly higher temperature throughout most of the stellar profile compared to the lower limit, on the order of $10^{7}\ \mathrm{K}$. Moreover, in contrast to the interior, the surface temperature is slightly lower by a comparable amount.
We observe a similar behavior at $t=10^{-1}$ yr, with a hotter interior and a steeper temperature gradient in the outer layers. However, at this point we begin to access the differences in the thermalization process between the limits considered. The configuration with $\bar{\Lambda}=4\mu$ appears to thermalize more rapidly than that with $\bar{\Lambda}=2\mu$, since it exhibits a more extended region within the star that is isothermal with the center. By the time both configurations reach $\mathrm{t}=1$ yr, the star becomes nearly isothermal, with a temperature difference between the center of the star ($r=0$) and the surface ($r=R_{\rm star}$) of the order of $10^{4}\ \mathrm{K}$ and $10^{5}\ \mathrm{K}$ for the lower and upper bounds, respectively. This difference decreases further to the order of $10^{0}$ K at $t=14$ yr for $\bar{\Lambda}=2\mu$ and $t=17$ yr for $\bar{\Lambda}=4\mu$.
@@
For quark stars featuring a hadronic crust, we compute the thermal evolution using two distinct numerical approaches, which we refer to as Grid Model A and Grid Model B. Grid Model A explicitly accounts for the presence of the gap between the quark core and the crust. In this case, when constructing the radial grid for the calculations, we place a grid point at the location of the gap by employing a sufficiently small radial step size ($dr$). Grid Model B, in contrast, neglects the gap entirely, initiating the crustal grid at the neutron drip density and continuing outward, down to a density of $\rho=10^{10}~\mathrm{g\,cm^{-3}}$.
While both approaches yield consistent results in the region of the cooling curve where observational data exist, they differ at early times: the explicit treatment of the gap affects the initial luminosity evolution and the early internal temperature profiles. Since this young epoch is not constrained by observations, and because resolving the narrow gap–of order several hundred femtometres–significantly increases computational cost, we adopt the second method when comparing our results with observational cooling data (including the comparison with the MIT bag model). Nevertheless, for completeness, we will explicitly show the differences between the two approaches by presenting the pQCD luminosity evolution and the corresponding internal temperature profiles obtained from both methods.
@@
Furthermore, the luminosity band of the crusted quark star configurations remains systematically broader than that of the bare case at all ages. As in the bare quark star scenario, the widening of the band is associated with the stellar age at which the luminosity curve exhibits its steepest decline. Since the upper and lower boundaries reach this point at distinct stellar ages, the luminosity spread is consequently stretched during this phase.
This behavior signals that the sensitivity to the choice of the renormalization scale becomes more pronounced in the cooling of quark stars with a hadronic crust.
@@
In Figure 8, we show the temperature profile evolution for the two modeling approaches for a quark star with a crust. The top panels correspond to the case where the thermal gap at the crust–core boundary is taken into account, whereas the bottom panels show the results when this gap is neglected. The most significant differences arise during the early evolutionary stages. At these times, as shown for $t=10^{-1}$ yr, $t=1$ yr and $t=10$ yr, the temperature profile exhibits a clear discontinuity between the crust (for densities $\rho\leq 4\times 10^{10}\ \mathrm{g}\ \mathrm{cm}^{-3}$) and the core ($\rho>4\times 10^{10}\ \mathrm{g}\ \mathrm{cm}^{-3}$) when we take the gap into account. The core rapidly becomes isothermal, while the crust retains higher temperatures, except in the outermost layers where surface photon emission plays an important role.
In contrast, when the thermal gap is not included, the transition between crust and core is smoother, and the core takes longer to reach an isothermal state. Despite these early-time differences in the internal temperature distribution, both models converge to nearly identical thermal profiles by $t=100$ yr. As a result, the differences between the predicted luminosities appear only for ages $\lesssim 100$ yr, as illustrated in Figure 9. Such ages, however, are significantly younger than those of any neutron stars in our observational sample. Consequently, the early–time discrepancy between the two grid models has no practical impact on the comparison with data.
Figure 9: Redshifted photon luminosity $L_{\infty}$ as a function of stellar age for a $M=1.4\ M_{\odot}$ quark star with a hadronic crust, comparing the cooling results obtained with Grid Model A and Grid Model B.
Moreover, the computational expense associated with Grid Model A is substantially higher. Accurately positioning a grid point inside the narrow crust–core interface requires resolving length scales on the order of femtometers, which makes the radial derivatives exceedingly steep and demands integration steps that are too small for the scale we need for this problem. Grid Model B avoids this difficulty while producing indistinguishable results for all relevant ages.
Beyond the variations in the cooling band behavior, the temperature profile evolution further highlights differences between the bare and crusted quark star models. In the presence of a crust, a pronounced temperature gradient develops, with the two regions cooling at different rates. This behavior resembles that of neutron star cooling, where efficient neutrino emission in the core causes it to cool more rapidly than outer layers, leading to a “cold wave” propagating outward toward the crust. This process delays the onset of an isothermal profile. Specifically, the time required for the temperature difference between the center ($r=0$) and the boundary ($r=R_{\rm boundary}$) to decrease to a value on the order of $10^{3}$ K is approximately $10^{4}$ yr for $\bar{\Lambda}=2\mu$ and $4\times 10^{4}$ yr for $\bar{\Lambda}=4\mu$.
@@ V Summary and outlook
Motivated by these findings, we extended our analysis to quark stars with a hadronic crust. The motivation arises from the observation that the uncertainty associated with the pQCD EoS is more significant at early stages of thermal evolution, when the star has not yet achieved internal isothermality and is primarily cooling via neutrino emission. Given the known role of the crust as a thermal insulator, which delays the establishment of an isothermal interior, we investigated how the uncertainty from the pQCD EoS influences this configuration. As expected, the temperature evolution in the crusted case is less steep than in the bare configuration, bringing the model predictions into closer agreement with the observational data, as illustrated in Figure 7. In particular, the luminosity band obtained for the crusted model encompasses the data point of PSR B2334+61, although the full extent of its reported uncertainties is not simultaneously accommodated.
In contrast to the bare case, the cooling band in the crusted configuration does not exhibit a significant narrowing throughout the timescale considered (up to $t=10^{5}$ yr), indicating a more persistent sensitivity to the EoS parameters and emphasizing the need to include equation-of-state uncertainties in cooling calculations.
We also computed the cooling curves obtained using the MIT bag model with $\mathcal{O}(\alpha_{s})$ corrections. For the bare configuration, the MIT bag model predictions remain within the uncertainty band associated with the variation of $\bar{\Lambda}$ in the pQCD equation of state, except during the very early stages of the evolution. In this initial phase, the MIT bag model curve briefly falls outside the band but enters it before the star reaches an age of $1$ yr. For quark stars with a hadronic crust, however, the behavior is reversed. In this case, the MIT bag model prediction begins within the pQCD uncertainty band but subsequently departs from it, yielding higher luminosities.
It is well established that color superconductivity can modify both the neutrino emissivity and the heat capacity of quark matter Shovkovy and Ellis (2002); Blaschke et al. (2006); Noda et al. (2013); Schmitt and Shternin (2018). A natural extension of the present study would therefore be to investigate the cooling of quark stars in different color-superconducting phases, while consistently incorporating the uncertainties from the pQCD equation of state. Another relevant direction is to refine the microphysical inputs by implementing up-to-date calculations of neutrino emissivity Schäfer and Schwenzer (2004); Adhya et al. (2013) and heat capacity in quark matter Gerhold et al. (2004).